# Realizable Continuous-Space Shields for Safe Reinforcement Learning

### v1

@@ 4 Continuous Shields
Example 1
For instance, consider the following non-markovian specification: if the robot is in region $r\in R$, then it cannot visit $r$ again in at least $i$ timesteps. Moreover, let dynamics of positions $(x,y)$ of the robot in step $k$ be precisely characterized by valuations of variables in step $k-1$. We can formally state this in:
$\square[(x=$ $\lhd\cos(\theta)\cdot\lhd d+x)\wedge(y=\lhd\sin(\theta)\cdot\lhd d+y)]$
$\implies\varphi\wedge\square$ $[\bigwedge_{r\in R}[(r_{b}^{x}\leq x\leq r_{u}^{x})\wedge(r_{b}^{y}\leq y\leq r_{u}^{y})]$
$\implies$ $\square_{[1,i]}\neg[(r_{b}^{x}\leq x\leq r_{u}^{x})\wedge(r_{b}^{y}\leq y\leq r_{u}^{y})]],$
where $\theta$ and $d$ correspond to the angle and distance decided by the system (based on environment observations that are modelled in the safety property $\varphi$). In natural language, the first line of the specification describes the dynamics of $(x,y)$, whereas the second and third line describe that if $(x,y)\in r$ then $(x,y)\notin r$ throughout $i$ timesteps. Note that this specification is syntactically encoded in a fragment of LTLt that is not decidable. However, we can rewrite it in the NCS fragment of LTLt with the formula $A\rightarrow G$, where $A$ are the following assumptions contraining the dynamics of the environment:
$\square[\bigwedge_{r\in R}\neg[(r^{x_{b}}\leq\cos(\theta)\cdot d+x\leq r^{x_{u}})$
$\wedge(r^{y_{b}}\leq\sin(\theta)\cdot d+y\leq r^{y_{u}})]$
$\rightarrow\ocircle\neg[(r^{x_{b}}\leq x\leq r^{x_{u}})\wedge(r^{y_{b}}\leq y\leq r^{y_{u}})]],$
which means that there is no situation in which $\theta$ and $d$ are in $r$ in timestep $k$ and $(x,y)$ of timestep $k+1$ is not in $i$). Moreover, $G$ are the following guarantees that encode what the system should enforce:
$\square[\bigwedge_{r\in R}[[(r^{x_{b}}\leq\cos(\theta)\cdot d+x\leq r^{x_{u}})\wedge$
$(r^{y_{b}}\leq\sin(\theta)\cdot d+y\leq r^{y_{u}})]$
$\rightarrow\square_{[1,i]}\neg[(r^{x_{b}}\leq\cos(\theta)\cdot d+x\leq r^{x_{u}})$
$\wedge(r^{y_{b}}\leq\sin(\theta)\cdot d+y\leq r^{y_{u}})]]],$
which means that, if $\theta$ and $d$ are in $r$ in timestep $k$, then throught $i$ timesteps, $\theta$ and $d$ must be chosen so that they are not in $r$).
@@
5 Case Study: Mapless Navigation
In this section, we present the main case study of this work, describing the environment we use, the safety challenges it presents, and the safety specifications we aim to guarantee with a proper shield over our RL agent.
Navigation with Reinforcement Learning
The goal of our agent is to navigate through an unknown environment to reach a target position while avoiding collisions with obstacles. This problem is particularly challenging from a safety perspective due to the partial observability of the environment. In every episode, we randomly set the position of the agent’s starting point, target destination, and the placement of obstacles. The agent can only perceive its immediate environment via a lidar scan, providing scalar distances to nearby obstacles in a finite set of directions. In addition to lidar, observations include the agent’s current position, orientation, and target location. We define the environment’s action space as the linear and angular velocity of the robot (we provide additional details about the training environment in Appendix B). The reward function for this environment incentivizes the agent to reach its target position by avoiding the obstacles and additionally provides a penalty to encourage the robot to reach the target with the minimum possible number of steps; more formally:
Figure 3: Average success rate and collision rate obtained during the DRL training process for 500 episodes (x-axis) without any shield applied (averaged over $5$ different random seeds). We reach an imperfect final collision rate of 1.2%, highlighting the need for a shield to prevent dangerous failures.
@@ 5 Case Study: Mapless Navigation
To train our agent, we employ Proximal Policy Optimization (PPO) (Schulman et al. 2017), which has been shown to be efficient on similar navigation tasks (Amir et al. 2023). For evaluation, we analyze the success and collision rates of trajectories executed by the agent. Fig.3 illustrates the PPO training process over time. Without any shielding, we see the agent achieve a success rate of 98.1% along with a complementarily low collision rate of 1.2%. However, the collision rate does not reach a constant value of zero, highlighting the persistent risk of unsafe behavior during deployment. This observation underscores the critical need for our proposed shielding approach, which aims to enforce robust safety guarantees and mitigate the remaining risks inherent in complex and dynamic environments.
Continuous Shield for Safe Navigation
@@
In this section, we delineate the safety requirements we aim to enforce and how we specify them using LTLt to ensure robust and reliable performance. We consider two key aspects: safety (i.e. avoiding collisions) and liveliness (i.e. avoiding repeated action sequences that can lead to a loop).
Requirements for Collision Avoidance
@@
In navigation tasks, ensuring collision avoidance is crucial under all circumstances. To achieve this, we encode specifications by leveraging the dynamics of the robot, made feasible by our framework’s ability to handle continuous spaces. Given an action $a=[a^{0},a^{1}]$, the robot first rotates by $a^{1}$ and then translates by $a^{0}$, where $a^{0}\in[-L^{0},L^{0}]$ and $a^{1}\in[-L^{1},L^{1}]$ and $L^{N}$ is a step size. Positive $a^{0}$ indicates forward motion and positive $a^{1}$ means a right turn. As the next pose of the robot is predictable through the dynamics, we can design requirements that strictly avoid collisions in the next timestep. Since robot rotation and translation are performed sequentially in the environment dynamics, requirements for each can be designed separately. Specifically, the red area in Fig. 4 (left) shows the robot’s trajectory when it makes a maximum right turn of $L^{1}$. To conservatively prevent collisions from right turns, we prohibit turning right if any lidar value $l^{i}$ is below a certain threshold $T^{i}$. These thresholds are precomputed based on the robot’s dynamics and maximum step size $L^{1}$ and they vary for each lidar (highlighted in blue). Formally, the specification for a right turn is: $\exists i$ s.t. $\,(l^{i}\leq T^{i})\rightarrow a^{1}\leq 0$. The same principle applies to left turns.
@@
For translation, we limit the maximum distance that we can translate based on the minimum distance of potential obstacles that any lidars sense in the direction of travel. Consider the $i^{th}$ lidar with value $l^{i}$ at an angular position $\theta^{i}$ relative to the robot’s horizontal line, measured clockwise from the left. Note that $\theta^{i}$ is known by construction as part of the system. After the robot rotates by $a^{1}$, the angular position w.r.t. the new horizontal line becomes $\hat{\theta^{i}}=\theta^{i}-a^{1}$. If this lidar reveals that the obstacle is in the path of forward translation, the translation $a^{0}$ cannot exceed the distance to the obstacle; otherwise, a collision will occur. Formally, the requirement for forward translation w.r.t. the $i^{th}$ lidar is: $l^{i}|\textrm{cos}{\theta^{i}}^{\prime}|\leq\frac{W}{2}\rightarrow a^{0}+H^{f}<l^{i}|\textrm{sin}{\theta^{i}}^{\prime}|$, where $W$ is the robot’s width and $H^{f}$ is the lidar’s forward offset. For forward translation, this requirement must hold for all lidars in front of the robot. Similar requirements apply for backward translation, with different signs, offsets, and considering lidars at the back.
@@
Table 2: Number of episodes where the shield returns unsat out of 500 test episodes with different $L^{Q}$ and $G^{A}$. Configurations verified as realizable are highlighted in green, otherwise in red. realizability checking can reveal the existence of situations in which specifications are unsatisfiable that empirical evaluation may fail to indicate.
Requirements for Liveness
While safety remains paramount in autonomous systems, ensuring collision avoidance, etc. alone can lead to overly conservative behavior, often characterized by repetitive or looping actions. The agent then can be particularly inefficient, wasting time and energy. To address this issue, we introduce a set of important liveness requirements. These prohibit repeating the same action in a state for a certain time window. This encourages the agent to explore new actions, potentially helping it escape from being stuck. We encode these requirements using a queue constructed from the shield state of length $L^{Q}$. Each queue element consists of the robot’s pose (x/y position and rotation) and actions, represented as $(x,y,r,a^{0},a^{1})$. To effectively check for repeated tuples of continuous values, we quantize the pose and action space into grids of $G^{P}$ and $G^{A}$ cells, respectively. In this way, the shield does not allow an action $(a^{0},a^{1})$ in a state $(x,y,r)$ if any of the previous tuples in the queue falls into the same cell. As the liveness requirements take stateful memory as an additional input, they are non-Markovian requirements and however, don’t affect the realizability check in our approach.
Realizability and Shielding Implementation
@@
We specify the above requirements in LTLt and provide them to our offline realizability check process to guarantee that these safety and liveness requirements are always satisfiable. Our realizability check process follows the implementation established in (Rodríguez and Sánchez 2023, 2024a, 2024b; Rodríguez, Gorostiaga, and Sánchez 2024). After realizability for our requirements is guaranteed, at test-time, we deploy our agent with an online shield to provide safe alternative actions when the agent suggests an action that violates our constraints. We present our LTLt requirements and the overall algorithm for deploying the online shield in the appendices.
Figure 5: An unshielded agent collides with obstacles and avoids them when a collision shield is applied. By introducing liveness requirements, the agent also avoids oscillations and reaches its target destination.
@@
No Shield Safety Shield Safety & Liveness Shield Optimizer Success Collision Success Collision Success Collision Success Collision Expert A 0.87 $\pm$ 0.05 0.03 $\pm$ 0.03 0.87 $\pm$ 0.04 0.00 $\pm$ 0.00 0.90 $\pm$ 0.02 0.00 $\pm$ 0.00 0.92 $\pm$ 0.02 0.00 $\pm$ 0.00 Expert B 0.88 $\pm$ 0.03 0.03 $\pm$ 0.02 0.87 $\pm$ 0.03 0.00 $\pm$ 0.00 0.90 $\pm$ 0.02 0.00 $\pm$ 0.00 0.90 $\pm$ 0.02 0.00 $\pm$ 0.00 Moderate A 0.77 $\pm$ 0.04 0.01 $\pm$ 0.01 0.76 $\pm$ 0.04 0.00 $\pm$ 0.00 0.79 $\pm$ 0.04 0.00 $\pm$ 0.00 0.80 $\pm$ 0.05 0.00 $\pm$ 0.00 Moderate B 0.85 $\pm$ 0.02 0.04 $\pm$ 0.02 0.85 $\pm$ 0.02 0.00 $\pm$ 0.00 0.87 $\pm$ 0.02 0.00 $\pm$ 0.00 0.87 $\pm$ 0.02 0.00 $\pm$ 0.00 Unsafe 0.22 $\pm$ 0.03 0.78 $\pm$ 0.03 0.40 $\pm$ 0.05 0.00 $\pm$ 0.00 0.47 $\pm$ 0.02 0.01* $\pm$ 0.01 0.69 $\pm$ 0.03 0.01* $\pm$ 0.00
@@
6 Experimental Results
In this section, we present the results of our experiments, demonstrating the effectiveness of our proposed realizability check and shielding approach.
Mapless Navigation Experiments
Realizability Check
A crucial aspect of our shield is that it guarantees a safe action in any state through a realizability check. Ensuring realizability by hand is challenging with complex requirements like ours, and statistics-driven safety checks are also unreliable.
In Tab. 2 we evaluate a shielded agent for 500 episodes with multiple different queue lengths ($L^{Q}$) and action grid sizes ($G^{A}$) without performing a realizability check beforehand. We then tally the number of times we encountered a unsat output from the shield due to no safe action being available. Finally, we verify the realizability of each shield specification and highlight the realizable configuration in green. We see that performing a realizability check can warn us about potentially unsatisfiable and thus unsafe specifications even when empirical evaluations do not indicate a problem. For instance, no unsat situations occurred for our shield with evaluating $(L^{Q},G^{A})=(100,30)$, however, our realizability check reveals that there still are potential situations where $(100,30)$ is not satisfiable. For subsequent experiments, we use one of the verified realizable configurations, $(L^{Q},G^{A})=(30,13)$.
Online Shielding
@@ 6 Experimental Results
The first column shows the success rate and collision rate when each policy is executed without any external shielding component. Although the success rates for the well-trained agents could be considered satisfactory, all agents made some collisions with an obstacle. The second column shows each policy’s performance with the collision avoidance shield added, clearly demonstrating the effectiveness of our shield, which reduces the number of collisions to zero. However, the shield alone leads to overly conservative behavior, resulting in the unsafe trajectories being converted into timeouts with oscillating behavior rather than successful episodes. Crucially, the third column shows the complete version of our shield (i.e., with both collision avoidance and liveness requirements), showing that the satisfaction of both properties allows the agent to recover from the conservative behavior and increase the success rate while ensuring the safety of the policy. Finally, the last column shows a small additional impact of optimizing the shield’s choice of safe action. Since we operate in a continuous space, our shield can employ an optimizer to return a safe action that minimizes distance to the original policy output. In this navigation domain, we minimize $|{\hat{a}^{1}}-a^{1}|$ where $a^{1}$ is the agent’s original angular velocity and ${\hat{a}^{1}}$ is the angular velocity of the shielded action. Fig. 5 illustrates a policy’s behavior both with and without a shield employed for collisions and liveness.
Particle World Experiments
To further demonstrate the generalizability and robustness of our shielding approach, we also experiment in a multi-agent Particle World environment (Mordatch and Abbeel 2017). Four agents are tasked with reaching target positions on the opposite side of the map while maintaining a safe distance from one another. Fig. 2 (Right) shows a screenshot of this environment where agents are represented as dots. The primary safety requirement in this environment is to ensure that the agents always keep a specified minimum distance between each other, illustrated as circles around the agents. The results, summarized in Table 4, demonstrate that our safety shield can be seamlessly applied to this new environment with a continuous state and action space. Notably, our shielding technique in Particle World successfully eliminates all violations of the safety requirements, effectively preventing any unsafe actions. This highlights the versatility and effectiveness of our shielding approach, showcasing its potential for broader applications across various continuous-space environments.

### v2

@@ 3.1 Non-Markovian Encoding in LTLt
where (i) $A$ (left) is an environment assumption that describes the dynamics of the scenario and allows to compute the next state given the current state and the action. Specifically, the operator $\lhd x$ represents a specific value that the variable x has taken in the previous timestep; and (ii) $G$ (right) are the guarantees by the system, in this case, stating that if $x\in r$ in a given timestep, then $x\notin r$ throughout $i$ timesteps; i.e., the region $r$ will not be re-visited for $i$ timesteps.
Example 3.4.
@@
$[\tilde{A}]\hskip 14.22636pt\square[\bigwedge_{r\in R}[(x+a)\in r]\rightarrow\ocircle[x\in R]],$
@@
An essential observation that makes this transformation possible is that we are (i) assuming to have access to a safety-relevant subset of the robot’s dynamic, often called safety-dynamics in the literature (Yang et al., 2023); and (ii) we encoded the requirement for a finite number of regions $r\in R$ and a finite horizon $i$. In our example, we can anticipate the constraints-relevant components of the next state given the action and the current state of the system. Thus, we can anticipate any unsafe action of the agent, and potentially override it (whenever the specification realizable).
Theorem 3.5.
[…]
Definition 3.6 (Anticipation Fragment).
[…]
Corollary 3.7.
If $\varphi$ belongs to the anticipation fragment, then the realizability of $\varphi$ is decidable.
4 Case Study: Mapless Navigation with Reinforcement Learning
@@
The agent’s goal is to navigate through an unknown environment to reach a target position while avoiding collisions with obstacles. This problem is particularly challenging from a safety perspective due to the partial observability of the environment. Each episode randomizes the agent’s starting point, target destination, and obstacle placements. The agent can only perceive its immediate environment via a lidar scan, providing scalar distances to nearby obstacles in a finite set of directions. In addition to lidar, observations include the agent’s current position, orientation, and target location. Actions consist of linear and angular velocities of the robot, and training is performed using Proximal Policy Optimization (PPO) (Schulman et al., 2017), which has been shown to be efficient on similar navigation tasks (Amir et al., 2023). Additional details on the training environment, reward function, and algorithm are provided in Appendix B. For evaluation, we analyze the success and collision rates of trajectories executed by the agent. Fig.3 illustrates the training process over time. Without any shielding, the agent achieves a success rate of 98.1% along with a complementarily low collision rate of 1.2%. However, the collision rate does not reach a constant value of zero, highlighting the persistent risk of unsafe behavior during deployment. This observation underscores the critical need for our proposed shielding approach, which aims to enforce robust safety guarantees and mitigate the remaining risks inherent in complex and dynamic environments.
4.1 Continuous Shield for Safe Navigation
@@
In this section, we delineate the safety requirements we aim to enforce and how we specify them using LTLt to ensure robust and reliable performance.
Markovian Requirements for Collision Avoidance.
@@
In navigation tasks, ensuring collision avoidance is crucial under all circumstances. To achieve this, we encode specifications by leveraging the robot’s dynamics, which our framework supports for continuous spaces. Given an action $a=[a^{0},a^{1}]$, the robot first rotates by $a^{1}$ and then translates by $a^{0}$, where $a^{0}\in[-L^{0},L^{0}]$ and $a^{1}\in[-L^{1},L^{1}]$ and $L^{N}$ is a step size. Positive $a^{0}$ indicates forward motion and positive $a^{1}$ means a right turn. As the next pose of the robot is predictable through the dynamics, we can design requirements that strictly avoid collisions in the next timestep. Since robot rotation and translation are performed sequentially in the environment dynamics, requirements for each can be designed separately. Specifically, the red area in Fig. 3 (left) shows the robot’s trajectory when it makes a maximum right turn of $L^{1}$. To conservatively prevent collisions from right turns, we prohibit turning right if any lidar value $l^{i}$ is below a certain threshold $T^{i}$. These thresholds are precomputed based on the robot’s dynamics and maximum step size $L^{1}$ and they vary for each lidar (highlighted in blue). Formally, the specification for a right turn is: $\exists i$ s.t. $\,(l^{i}\leq T^{i})\rightarrow a^{1}\leq 0$. The same principle applies to left turns. For translation, we limit the maximum distance that we can translate based on the minimum distance of potential obstacles that any lidars sense in the direction of travel. Consider the $i^{th}$ lidar with value $l^{i}$ at an angular position $\theta^{i}$ relative to the robot’s horizontal line, measured clockwise from the left. Note that $\theta^{i}$ is known by construction as part of the system. After the robot rotates by $a^{1}$, the angular position w.r.t. the new horizontal line becomes $\hat{\theta^{i}}=\theta^{i}-a^{1}$. If this lidar reveals that the obstacle is in the path of forward translation, the translation $a^{0}$ cannot exceed the distance to the obstacle; otherwise, a collision will occur. Formally, the requirement for forward translation w.r.t. the $i^{th}$ lidar is: $l^{i}|\textrm{cos}{\theta^{i}}^{\prime}|\leq\frac{W}{2}\rightarrow a^{0}+H^{f}<l^{i}|\textrm{sin}{\theta^{i}}^{\prime}|$, where $W$ is the robot’s width and $H^{f}$ is the lidar’s forward offset. For forward translation, this requirement must hold for all lidars in front of the robot. Similar requirements apply for backward translation, with different signs, offsets, and considering lidars at the back.
Figure 2: Average success rate and collision rate obtained during the DRL training process for 500 episodes (x-axis) without any shield applied (averaged over $5$ different random seeds).
@@
Non-Markovian Requirements for Loop Avoidance.
While safety remains paramount in autonomous systems, ensuring collision avoidance, etc. alone can lead to overly conservative behavior, often characterized by repetitive or looping actions. The agent then can be particularly inefficient, wasting time and energy. To address this issue, we introduce a set of important non-Markovian requirements. These prohibit repeating the same action in a state for a certain time window. This encourages the agent to explore new actions, potentially helping it escape from being stuck. We encode these requirements using a queue constructed from the shield state of length $L^{Q}$. Each queue element consists of the robot’s pose (x/y position and rotation) and actions, represented as $(x,y,r,a^{0},a^{1})$. To effectively check for repeated tuples of continuous values, we quantize the pose and action space into grids of $G^{P}$ and $G^{A}$ cells, respectively. In this way, the shield does not allow an action $(a^{0},a^{1})$ in a state $(x,y,r)$ if any of the previous tuples in the queue falls into the same cell. As these requirements take stateful memory as an additional input, they are non-Markovian requirements and however, don’t affect the realizability check in our approach.
4.2 Realizability and Shielding Implementation
@@
Table 1: Number of episodes where the shield returns unsat out of 500 tests, with realizable configurations in green and others in red. realizability checking identifies unsatisfiable specifications that empirical evaluation might overlook.
@@
We specify the above requirements in LTLt and provide them to the offline realizability check process to guarantee that these requirements are always satisfiable. Our realizability check process follows the implementation established in Rodríguez and Sánchez (2023, 2024a, 2024b); Rodríguez et al. (2024). After realizability for our requirements is guaranteed, at test-time, we deploy our agent with an online shield to provide safe alternative actions when the agent suggests an action that violates our constraints. The overall algorithm for deploying the online shield is presented in Appendix C.
5 Experimental Results
Realizability Check.
A crucial aspect of our shield is that it guarantees a safe action in any state through a realizability check. Ensuring realizability by hand is challenging with complex requirements like ours, and statistics-driven safety checks are also unreliable. In Table 1, we evaluate a shielded agent for 500 episodes with multiple different queue lengths ($L^{Q}$) and action grid sizes ($G^{A}$) without performing a realizability check beforehand. We then tally the number of times we encountered a unsat output from the shield due to no safe action being available. Finally, we verify the realizability of each shield specification and highlight the realizable configuration in green. We see that performing a realizability check can warn us about potentially unsatisfiable and thus unsafe specifications even when empirical evaluations do not indicate a problem. For instance, no unsat situations occurred for our shield with evaluating $(L^{Q},G^{A})=(100,30)$, however, our realizability check reveals that there still are potential situations where $(100,30)$ is not satisfiable. For subsequent experiments, we use one of the verified realizable configurations, $(L^{Q},G^{A})=(30,13)$.
Online Shielding.
@@ 5 Experimental Results
No Shield Collision Shield Collision & Loop Shield Optimizer Success Collision Success Collision Success Collision Success Collision Expert A 0.87 $\pm$ 0.05 0.03 $\pm$ 0.03 0.87 $\pm$ 0.04 0.00 $\pm$ 0.00 0.90 $\pm$ 0.02 0.00 $\pm$ 0.00 0.92 $\pm$ 0.02 0.00 $\pm$ 0.00 Expert B 0.88 $\pm$ 0.03 0.03 $\pm$ 0.02 0.87 $\pm$ 0.03 0.00 $\pm$ 0.00 0.90 $\pm$ 0.02 0.00 $\pm$ 0.00 0.90 $\pm$ 0.02 0.00 $\pm$ 0.00 Moderate A 0.77 $\pm$ 0.04 0.01 $\pm$ 0.01 0.76 $\pm$ 0.04 0.00 $\pm$ 0.00 0.79 $\pm$ 0.04 0.00 $\pm$ 0.00 0.80 $\pm$ 0.05 0.00 $\pm$ 0.00 Moderate B 0.85 $\pm$ 0.02 0.04 $\pm$ 0.02 0.85 $\pm$ 0.02 0.00 $\pm$ 0.00 0.87 $\pm$ 0.02 0.00 $\pm$ 0.00 0.87 $\pm$ 0.02 0.00 $\pm$ 0.00 Unsafe 0.22 $\pm$ 0.03 0.78 $\pm$ 0.03 0.40 $\pm$ 0.05 0.00 $\pm$ 0.00 0.47 $\pm$ 0.02 0.01* $\pm$ 0.01 0.69 $\pm$ 0.03 0.01* $\pm$ 0.00
@@
The first column shows the success rate and collision rate when each policy is executed without any external shielding component. Although the success rates for the well-trained agents could be considered satisfactory, all agents made some collisions with an obstacle. The second column shows each policy’s performance with the collision avoidance shield added, clearly demonstrating the effectiveness of our shield, which reduces the number of collisions to zero. However, the shield alone leads to overly conservative behavior, resulting in the unsafe trajectories being converted into timeouts with oscillating behavior rather than successful episodes. Crucially, the third column shows the complete version of our shield (i.e., with collision and loop avoidance requirements), showing that the satisfaction of both properties allows the agent to recover from the conservative behavior and increase the success rate while ensuring the safety of the policy. Finally, the last column shows a small additional impact of optimizing the shield’s choice of safe action. Since we operate in a continuous space, our shield can employ an optimizer to return a safe action that minimizes distance to the original policy output. In this navigation domain, we minimize $|{\hat{a}^{1}}-a^{1}|$ where $a^{1}$ is the agent’s original angular velocity and ${\hat{a}^{1}}$ is the angular velocity of the shielded action. Fig. 4 illustrates a policy’s behavior both with and without a shield employed for collisions and loop. In the literature, approaches for enhancing safety in deep reinforcement learning provide only empirical results that satisfy requirements in expectation, without formal guarantees (Ray et al., 2019; Corsi et al., 2022). While these methods achieve near-optimal safety on average, they fail to fully eliminate violations, underscoring the need for a shield, as discussed in Appendix F.
Figure 4: An unshielded agent collides with obstacles and avoids them when a collision shield is applied. By introducing non-Markovian requirements, the agent also avoids oscillations.
@@
Particle World Experiments
To further demonstrate the generalizability and robustness of our shielding approach, we also experiment in a multi-agent Particle World environment Mordatch and Abbeel (2017). Four agents are tasked with reaching target positions on the opposite side of the map while maintaining a safe distance from one another (a screenshot of this environment can be found in Fig. 5 of the supplementary materials). The primary safety requirement in this environment is to ensure that the agents always keep a specified minimum distance between each other, illustrated as circles around the agents. The results, summarized in Table 3, demonstrate that our safety shield can be seamlessly applied to this new environment with a continuous state and action space. Notably, our shielding technique in Particle World successfully eliminates all violations of the safety requirements, effectively preventing any unsafe actions. This highlights the versatility and effectiveness of our shielding approach, showcasing its potential for broader applications across various continuous-space environments.